### v1

# Crossing Numbers of Beyond Planar Graphs Re-revisited: A Framework Approach

Abstract
Beyond planarity concepts (prominent examples include $k$-planarity or fan-planarity) apply certain restrictions on the allowed patterns of crossings in drawings. It is natural to ask, how much the number of crossings may increase over the traditional (unrestricted) crossing number. Previous approaches to bound such ratios, e.g. [CKMV19, BPS21], require very specialized constructions and arguments for each considered beyond planarity concept, and mostly only yield asymptotically non-tight bounds. We propose a very general proof framework that allows us to obtain asymptotically tight bounds, and where the concept-specific parts of the proof typically boil down to a couple of lines. We show the strength of our approach by giving improved or first bounds for several beyond planarity concepts.
keywords:
Beyond planarity, crossing number, crossing ratio, proof framework
category:
[…]
1 Introduction
Throughout this paper, we only consider simple graphs, i.e., no self-loops or multi-edges. Given a graph $G$, let $n$ and $m$ denote its number of vertices $V(G)$ and edges $E(G)$, respectively. A drawing $\mathcal{D}$ of $G$ is a mapping of $V(G)$ to distinct points in $\mathbb{R}^{2}$, and $E(G)$ to curves connecting the respective end points. Such a curve must not contain any vertex point other than its end points. When two edge curves intersect on an internal point $x$, this is a crossing of these two edges at $x$. The crossing number $\mathrm{cr}(G)$ is the smallest number of crossings over all drawings. To achieve it, we can safely assume simple drawings, i.e., no three edges cross at a common point, no edge crosses itself or adjacent edges, and no pair of edges crosses multiple times. A graph is planar if it allows a crossing-free drawing. In the last decades, several beyond planarity concepts have been established that generalize planar graphs in that certain special crossing patters are allowed or forbidden, see, e.g., [SurveyBeyondplanarity, beyondBook] for overviews. A prominent example is $k$-planarity, where any edge may be crossed at most $k$ times.
We are interested in the minimum number of crossings within any beyond-planarity restricted drawing. Intuitively speaking, a beyond planarity concept is a way to (try to) formalize our understanding of what aspects constitute a readable drawing. Still, within these restrictions it is natural to ask for a drawing with the least number of crossings. Alternatively, we can ask how much it “costs” in terms of crossings to follow a certain drawing paradigm.
Formally, let $\mathfrak{B}(k)$ denote any beyond planarity concept, where we omit the parameter $k$ for parameterless beyond planarities. Then, $\mathcal{G}_{\mathfrak{B}(k)}$ ($\mathcal{G}_{\mathfrak{B}(k)}(n)$) denotes the set of graphs (on $n$ vertices, respectively) allowing a $\mathfrak{B}(k)$-drawing. The $\mathfrak{B}(k)$-crossing number $\mathrm{cr}_{\mathfrak{B}(k)}(G)$ for a graph $G\in\mathcal{G}_{\mathfrak{B}(k)}$ is the least number of crossings over all $\mathfrak{B}(k)$-drawings of $G$. The crossing ratio $\varrho_{\mathfrak{B}(k)}(n)$ of $\mathfrak{B}(k)$ is the largest attainable ratio between the $\mathfrak{B}(k)$-crossing number and the (normal) crossing number, over all $n$-vertex $\mathfrak{B}(k)$-graphs:
$\varrho_{\mathfrak{B}(k)}(n)\coloneqq\sup_{G\in\mathcal{G}_{\mathfrak{B}(k)}(n)}\frac{\mathrm{cr}_{\mathfrak{B}(k)}(G)}{\mathrm{cr}(G)}.$
Crossing ratios of beyond planar concepts have for the first time been explicitly considered in [CKMV19], where linear lower bounds $\Omega(n)$ where established for 1-planar, $k$-quasi-planar, and (weakly) fan-planar graphs. In all but the first case, the upper bounds where (at least) quadratic in $n$, and it was conjectured in [CKMV19] that the real lower bounds should rather be $\Omega(n^{2})$. Later, [BPS21] generalized the results to $k$-planarity and seven more beyond planarity concepts. However, none of the provided bounds were tight except for planarly-connected and straight-line RAC. In both publications, each planarity measure needs a very specific construction and intricate specialized arguments to prove the lower bounds (which tend to always be the most complicated part in crossing number proofs). Furthermore, both only consider simple drawings, when establishing upper bounds on the crossing ratio. Simple and non-simple crossing ratios may differ, if there are $\mathfrak{B}(k)$-graphs that only allow non-simple $\mathfrak{B}(k)$-drawings. Our bounds below work with and without the simplicity assumption.
Our contribution.
In this paper, we propose a general framework that, when applicable, simplifies crossing ratio proofs down to only a couple of $\mathfrak{B}(k)$-specific lines (10–20 lines for the lower bound). In contrast to previous schemes, this framework is able to prove tight bounds for all considered beyond planarities, and our proofs further show that the crossing ratio is already achieved between any two subsequent parameterizations $k$ and $k+1$, in all considered parametrized concepts. Table 1 summarizes our results. A key idea lies in the simplification of the necessary lower bound arguments, by turning them into counting arguments over a set of (sufficiently) disjoint Kuratowski-subdivisions (see Section 2). In particular, this allows us to avoid any intricate discussions about alternative drawing possibilities, the simplicity of drawings, etc. We present the proof framework in Section 3, and showcase its versatility and strength for various beyond planarity concepts in LABEL:sec:use.
Table 1: Overview on our results on the crossing ratios of beyond planarities for $n$-vertex graphs. Formally, [CKMV19] considers weakly fan-planar, but their proof works for all 4 variants. The column “$\overline{\varrho}$?” denotes whether we also show the (same) rectilinear crossing ratios, see LABEL:sec:rectilinear. The -marked previous upper bounds need no explicit discussion of non-simple drawings by the nature of the beyond planarity concept. All our bounds work with and without the simplicity assumption.
beyond planarity previous best our results Section $\overline{\varrho}$?
$k$-planar $\Omega(n/k)\cap\mathcal{O}(k\sqrt{k}n)$ [BPS21] $\Theta(n)$ LABEL:sec:kpl
$k$-vertex-planar $\Theta(n)$ LABEL:sec:kvp
IC-planar $\Theta(n)$ LABEL:sec:ic
NIC-planar $\Theta(n)$ LABEL:sec:nic
NNIC-planar $\Theta(n^{2})$ LABEL:sec:nnic
$k$-fan-crossing-free $\Omega(n^{2}/k^{3})\cap\mathcal{O}(k^{2}n^{2})$ [BPS21] $\Theta(n^{2}/k)$ LABEL:sec:nnic
straight-line RAC $\Theta(n^{2})$ [BPS21] (direct corollary) LABEL:sec:nnic
adjacency-crossingfan-crossingweakly fan-planarstrongly $\begin{array}[]{@{}l}\text{adjacency-crossing}\\ \text{fan-crossing}\\ \text{weakly fan-planar}\\ \text{strongly fan-planar}\\ \end{array}\left.\begin{array}[]{@{}l}\\ \\ \\ \\ \end{array}\right\}$ $\Omega(n)\cap\mathcal{O}(n^{2})$ [CKMV19] $\Theta(n^{2})$ LABEL:sec:fp $\times$
$k$-edge-crossing $\Theta(k)$ LABEL:sec:ecr
$k$-gap-planar $\Omega(n/k^{3})\cap\mathcal{O}(k\sqrt{k}n)$ [BPS21] $\Theta(n/k)$ LABEL:sec:kgappl
$k$-apex-planar $\Omega(n/k)\cap\mathcal{O}(k^{2}n^{2}+k^{4})$ [BPS21] $\Theta(n^{2}/k)$ LABEL:sec:kapex
skewness-$k$ $\Omega(n/k)\cap\mathcal{O}(kn+k^{2})$ [BPS21] $\Theta(n)$ LABEL:sec:kapex
2 Preliminaries
Kuratowski subdivisions.
A graph $G$ is planar if and only if it does not contain a $K_{5}$ or $K_{3,3}$ subdivision (summarily called Kuratowski subdivisions[kuratowski1930probleme]. Let $K$ be a $K_{3,3}$ subdivision in $G$. The six degree-3 vertices of $K$ are the Kuratowski nodes of $K$, and the paths between them (not containing other Kuratowski nodes) its Kuratowski paths; by replacing each Kuratowski path by a single edge, one obtains $K_{3,3}$. Two Kuratowski paths are adjacent, if they share a common Kuratowski node. Clearly, we have $\mathrm{cr}(G)\geq 1$, and any drawing $\mathcal{D}$ of $G$ (even a non-simple one) contains at least one crossing $x$ between two edges from non-adjacent Kuratowski paths. We say that $x$ covers $K$. Every Kuratowski subdivision in $G$ has to be covered in $\mathcal{D}$; a single crossing may cover several such subdivisions.
Upper bounds.
The crossing lemma [CrossingLemmaPaper] states that any graph $G$ with $n$ vertices and $m>4n$ edges has $\mathrm{cr}(G)\in\Omega(\frac{m^{3}}{n^{2}})$. We may split $\mathcal{G}_{\mathfrak{B}(k)}(n)$ into graphs $\mathcal{G}^{\mathrm{dense}}_{\mathfrak{B}(k)}(n)\coloneqq\{G\in\mathcal{G}_{\mathfrak{B}(k)}(n)\mid m>4n\}$ that are sufficiently dense for the crossing lemma, and $\mathcal{G}^{\mathrm{sparse}}_{\mathfrak{B}(k)}(n)\coloneqq\mathcal{G}_{\mathfrak{B}(k)}(n)\setminus\mathcal{G}^{\mathrm{dense}}_{\mathfrak{B}(k)}(n)$. It will later turn out that we attain the largest ratios on graphs of the latter subset. Let $\varrho_{\mathfrak{B}(k)}^{\mathrm{sparse}}(n)\coloneqq\sup_{G\in\mathcal{G}^{\mathrm{sparse}}_{\mathfrak{B}(k)}(n)}\frac{\mathrm{cr}_{\mathfrak{B}(k)}(G)}{\mathrm{cr}(G)}$. Then
$\varrho_{\mathfrak{B}(k)}(n)\leq\max\left\{\varrho_{\mathfrak{B}(k)}^{\mathrm{sparse}}(n)\ ,\ \sup_{G\in\mathcal{G}^{\mathrm{dense}}_{\mathfrak{B}(k)}(n)}\frac{\mathrm{cr}_{\mathfrak{B}(k)}(G)}{\mathrm{cr}(G)}\right\},$
where we want to use the crossing lemma to bound the second term in the maximum. {observation} We have:
  1. (a)
    If $\mathrm{cr}_{\mathfrak{B}(k)}(G)\in\mathcal{O}(m^{2})$, then $\varrho_{\mathfrak{B}(k)}(n)\in{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\mathcal{O}(\varrho_{\mathfrak{B}(k)}^{\mathrm{sparse}}(n)+m^{2}\cdot\frac{n^{2}}{m^{3}})\subseteq}\;\mathcal{O}(\varrho_{\mathfrak{B}(k)}^{\mathrm{sparse}}(n)+n).$
  2. (b)
    If $\mathrm{cr}_{\mathfrak{B}(k)}(G)\in\mathcal{O}(km)$, then $\varrho_{\mathfrak{B}(k)}(n)\in{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\mathcal{O}(\varrho_{\mathfrak{B}(k)}^{\mathrm{sparse}}(n)+km\cdot\frac{n^{2}}{m^{3}})\subseteq}\;\mathcal{O}(\varrho_{\mathfrak{B}(k)}^{\mathrm{sparse}}(n)+k).$
All simple drawings satisfy the prerequisite of Section 2(a). Furthermore, in the context of connected sparse graphs in $\mathcal{G}^{\mathrm{sparse}}_{\mathfrak{B}(k)}(n)$ (and thus for $\varrho_{\mathfrak{B}(k)}^{\mathrm{sparse}}(n)$), we know $m\in\Theta(n)$.
1-jumping.
The crossing ratio $\varrho_{\mathfrak{B}(k)}(n)$ relates the $\mathfrak{B}(k)$-restricted crossing number to the normal one. It is natural to ask, how small a parameter $\Delta\geq 1$ can be chosen such that $\varrho_{\mathfrak{B}(k)}^{\Delta}(n)\coloneqq\sup_{G\in\mathcal{G}_{\mathfrak{B}(k)}(n)}\frac{\mathrm{cr}_{\mathfrak{B}(k)}(G)}{\mathrm{cr}_{\mathfrak{B}(k+\Delta)}(G)}$ attains the same asymptotic bound. We say $\varrho_{\mathfrak{B}(k)}(n)$ is $\Delta$-jumping if $\varrho_{\mathfrak{B}(k)}^{\Delta}(n)\in\Omega(\varrho_{\mathfrak{B}(k)}(n))$. In LABEL:sec:use, we will always observe 1-jumping as our upper bounds on the normal crossing numbers are attained by $\mathfrak{B}(k+1)$-drawings.
3 Framework for Proving Lower Bounds on Crossing Ratios

### v2

@@
 A prominent example is $k$-planarity, where any edge may be crossed at most $k$ times.
-We are interested in the minimum number of crossings within any beyond-planarity restricted drawing.
+We are interested in the minimum number of crossings within any beyond planarity restricted drawing.
 Intuitively speaking, a beyond planarity concept is a way to (try to) formalize our understanding of what aspects constitute a readable drawing.
@@
 Alternatively, we can ask how much it “costs” in terms of crossings to follow a certain drawing paradigm.
-Formally, let $\mathfrak{B}(k)$ denote any beyond planarity concept, where we omit the parameter $k$ for parameterless beyond planarities.
+Formally, let $\mathfrak{B}(k)$ denote any beyond planarity concept, where we omit the parameter $k$ for parameterless beyond planarity concepts.
 Then, $\mathcal{G}_{\mathfrak{B}(k)}$ ($\mathcal{G}_{\mathfrak{B}(k)}(n)$) denotes the set of graphs (on $n$ vertices, respectively) allowing a $\mathfrak{B}(k)$-drawing.
@@
 $\varrho_{\mathfrak{B}(k)}(n)\coloneqq\sup_{G\in\mathcal{G}_{\mathfrak{B}(k)}(n)}\frac{\mathrm{cr}_{\mathfrak{B}(k)}(G)}{\mathrm{cr}(G)}.$
-Crossing ratios of beyond planar concepts have for the first time been explicitly considered in [CKMV19], where linear lower bounds $\Omega(n)$ where established for 1-planar, $k$-quasi-planar, and (weakly) fan-planar graphs.
+Crossing ratios of beyond planarity concepts have for the first time been explicitly considered in [CKMV19a, CKMV19], where linear lower bounds $\Omega(n)$ where established for 1-planar, $k$-quasi-planar, and (weakly) fan-planar graphs.
 In all but the first case, the upper bounds where (at least) quadratic in $n$, and it was conjectured in [CKMV19] that the real lower bounds should rather be $\Omega(n^{2})$.
@@
 In this paper, we propose a general framework that, when applicable, simplifies crossing ratio proofs down to only a couple of $\mathfrak{B}(k)$-specific lines (10–20 lines for the lower bound).
-In contrast to previous schemes, this framework is able to prove tight bounds for all considered beyond planarities, and our proofs further show that the crossing ratio is already achieved between any two subsequent parameterizations $k$ and $k+1$, in all considered parametrized concepts.
+In contrast to previous schemes, this framework is able to prove asymptotically tight bounds for all considered beyond planarity concepts, and our proofs further show that the crossing ratio is already achieved between any two subsequent parameterizations $k$ and $k+1$, in all considered parametrized concepts.
 Table 1 summarizes our results.
@@
 We present the proof framework in Section 3, and showcase its versatility and strength for various beyond planarity concepts in LABEL:sec:use.
-Table 1: Overview on our results on the crossing ratios of beyond planarities for $n$-vertex graphs.
+Table 1: Overview on our results on the crossing ratios of beyond planarity concepts for $n$-vertex graphs.
 Formally, [CKMV19] considers weakly fan-planar, but their proof works for all 4 variants.
-The column “$\overline{\varrho}$?” denotes whether we also show the (same) rectilinear crossing ratios, see LABEL:sec:rectilinear.
+The column “$\overline{\varrho}$?” denotes whether we also show the (same) rectilinear crossing ratios in LABEL:sec:rectilinear.
 The -marked previous upper bounds need no explicit discussion of non-simple drawings by the nature of the beyond planarity concept.
-All our bounds work with and without the simplicity assumption.
-beyond planarity previous best our results Section $\overline{\varrho}$?
+All our bounds, except the upper bounds for adjacency-crossing and fan-crossing, work with and without the simplicity assumption.
+beyond planarity concept previous best our results Section $\overline{\varrho}$?
 $k$-planar $\Omega(n/k)\cap\mathcal{O}(k\sqrt{k}n)$ [BPS21] $\Theta(n)$ LABEL:sec:kpl
@@
 $k$-gap-planar $\Omega(n/k^{3})\cap\mathcal{O}(k\sqrt{k}n)$ [BPS21] $\Theta(n/k)$ LABEL:sec:kgappl
-$k$-apex-planar $\Omega(n/k)\cap\mathcal{O}(k^{2}n^{2}+k^{4})$ [BPS21] $\Theta(n^{2}/k)$ LABEL:sec:kapex
-skewness-$k$ $\Omega(n/k)\cap\mathcal{O}(kn+k^{2})$ [BPS21] $\Theta(n)$ LABEL:sec:kapex
+$k$-apex-planar $\Omega(n/k)\cap\mathcal{O}(k^{2}n^{2})$ [BPS21] $\Theta(n^{2}/k)$ LABEL:sec:kapex
+skewness-$k$ $\Omega(n/k)\cap\mathcal{O}(kn)$ [BPS21] $\Theta(n)$ LABEL:sec:kapex
 2 Preliminaries
@@
 Then
-$\varrho_{\mathfrak{B}(k)}(n)\leq\max\left\{\varrho_{\mathfrak{B}(k)}^{\mathrm{sparse}}(n)\ ,\ \sup_{G\in\mathcal{G}^{\mathrm{dense}}_{\mathfrak{B}(k)}(n)}\frac{\mathrm{cr}_{\mathfrak{B}(k)}(G)}{\mathrm{cr}(G)}\right\},$
+$\varrho_{\mathfrak{B}(k)}(n)=\max\left\{\varrho_{\mathfrak{B}(k)}^{\mathrm{sparse}}(n)\ ,\ \sup_{G\in\mathcal{G}^{\mathrm{dense}}_{\mathfrak{B}(k)}(n)}\frac{\mathrm{cr}_{\mathfrak{B}(k)}(G)}{\mathrm{cr}(G)}\right\},$
 where we want to use the crossing lemma to bound the second term in the maximum. {observation} We have:
@@
   2. (b)
-    If $\mathrm{cr}_{\mathfrak{B}(k)}(G)\in\mathcal{O}(km)$, then $\varrho_{\mathfrak{B}(k)}(n)\in{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\mathcal{O}(\varrho_{\mathfrak{B}(k)}^{\mathrm{sparse}}(n)+km\cdot\frac{n^{2}}{m^{3}})\subseteq}\;\mathcal{O}(\varrho_{\mathfrak{B}(k)}^{\mathrm{sparse}}(n)+k).$
+    If $\mathrm{cr}_{\mathfrak{B}(k)}(G)\in\mathcal{O}(mk)$, then $\varrho_{\mathfrak{B}(k)}(n)\in{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\mathcal{O}(\varrho_{\mathfrak{B}(k)}^{\mathrm{sparse}}(n)+mk\cdot\frac{n^{2}}{m^{3}})\subseteq}\;\mathcal{O}(\varrho_{\mathfrak{B}(k)}^{\mathrm{sparse}}(n)+k).$
 All simple drawings satisfy the prerequisite of Section 2(a).
@@
 We say $\varrho_{\mathfrak{B}(k)}(n)$ is $\Delta$-jumping if $\varrho_{\mathfrak{B}(k)}^{\Delta}(n)\in\Omega(\varrho_{\mathfrak{B}(k)}(n))$.
-In LABEL:sec:use, we will always observe 1-jumping as our upper bounds on the normal crossing numbers are attained by $\mathfrak{B}(k+1)$-drawings.
+For all concepts in LABEL:sec:use except $k$-gap-planarity (LABEL:sec:kgappl), we will observe 1-jumping as our upper bounds on the normal crossing numbers are already attained by $\mathfrak{B}(k+1)$-drawings.
 3 Framework for Proving Lower Bounds on Crossing Ratios